OpenStudy (anonymous):

(a + 2b)^3 = a^3 + 6a^2b +... Use the Binomial Theorem to complete the binomial expansion.; enter only the next two terms. Use the ^ key for exponents; example: 25x^2y+3y^2

OpenStudy (anonymous):

(x + y)^3 = x^3 + 3(x^2)y + 3(x)y^2 + y^3 Here x=a, and y=2b. Substitute and simplify. Show work.